Circular No. 6157

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1995J IN NGC 4512 W. Johnson, Anza, CA, reports his discovery of an apparent supernova

located ∼ 30′′ east and 30′′ north of the center of NGC 4512 (α = 12^{h}32^{m}.6, δ = +63^{o}54′, equinox 2000.0); the unfiltered CCD magnitude of SN 1995J on the two discovery frames, taken on Apr. 4.45 and 4.46 UT, is ∼ 16.5. Nothing appears at this position on the CCD Atlas of Deep Sky Objects (Buil and Thouvenot 1991, Sky Publ.), though there is a foreground star of mag ∼ 16 located ∼ 10′′ west and 25′′ south of the galaxy’s center. P. Garnavich, P. Challis, and R. Kirshner, Center for Astrophysics, report that an image taken by R. Jansen at the 1.2-m telescope on Mt. Hopkins on Apr. 7.2 clearly shows a new star (Johnson R magnitude 16.4 ± 0.1) at the position reported by Johnson; no object appears on the Palomar sky survey prints at that position.

AL COMAE BERENICES J. Patterson, Department of Astronomy, Columbia University, com-

municates: “Photometry of AL Com on Apr. 5/6 by D. Harvey, Center for Basement Astrophysics-West, reveals periodic waves of period 81.8 ± 1.4 min, with substantial power also at the first and second harmonics. These are likely to be some version of ‘superhumps’. The amplitude was 0.1 mag, but may grow in the days to come. The waveform was highly variable and suggested the possibility of multiple (noncommensurate) periods. Finding the correct periodic structure in these cases depends crucially on amassing a dense data stream, and we eagerly seek collaboration with observers who can contribute time-series photometry of this star (V -band or unfiltered). I can be reached at telephone 212-854-3276 (U.S.A.) or via e-mail

at jop@carmen.phys.columbia.edu.”

J. A. DeYoung, U.S. Naval Observatory, reports: “Time-resolved Kron-Cousins I-band photometry of AL Com with the 0.61-m USNO reflector in Washington, DC, on Apr. 6.010–6.094 and 7.016–7.158 UT shows a modulation consistent with a period of 81.2 min. A linear least-squares fit to the five available minima gives the following HJD ephemeris for minimum phase: 2449813.527 (±0.002) + 0.0564 (±0.0001)E. The peak-to-peak amplitudes on Apr. 6 and 7 were 0.11 ± 0.02 and 0.06 ± 0.02 mag, respectively. In addition, there is phase-repeatable structure in the light curves.”

1995 April 7 Daniel W. E. Green